# Decentralized Domain Generalization with Style Sharing: Formal Model and Convergence Analysis

### v3

@@ IV-B1 Global Loss
where $F_{i}(\theta,\Psi_{\theta,i}(X))$ is the local loss function of each device $i\in\mathcal{M}$ evaluated with model parameters $\theta$ and combined batch $X$ across devices, in which $\Psi_{\theta,i}(X)\subset\Psi_{\theta}(X)$ is the set of styles from neighbors of device $i$. Note that in Eq. (19), if instead of $\Psi_{\theta,i}(X)$ we provide the full style information of $\Psi_{\theta}(X)$ to each individual device, we recover the loss function for DG in star-topology FL, as considered in StableFDG [13].
@@ IV-B2 Local Loss
$F_{i}(\theta,\Psi)\hskip-2.84526pt=\hskip-11.38109pt\sum_{\pi\in\mathcal{P}(L)}{\hskip-5.69054pt\left(\prod_{\ell\in\pi}{p_{\ell}}\hskip-2.84526pt\right)\hskip-2.84526pt\mathbb{E}_{\hskip-8.53581pt\begin{subarray}{c}(x,y)\sim\mathbb{p}_{D_{i}}\\ \{j_{\ell}\in\mathcal{M}\setminus\{i\}\}_{\ell\in\pi}\\ \{\mathcal{R}_{\ell}\sim\mathbb{p}_{\mathcal{R}}\}_{\ell\in\pi}\end{subarray}\hskip-11.38109pt}{\left[\mathscr{L}{(\tilde{h}_{\theta,\{\psi_{j_{\ell}},\mathcal{R}_{\ell}\}_{l\in\pi}}(x),y)}\right]}}\hskip-1.42262pt,\vskip-2.84526pt$ (20)
in which the expectation is taken over the set of random variables $\mathcal{R}_{\ell}=\left\{\epsilon_{\mu,\ell},\epsilon_{\sigma,\ell},\lambda_{\ell},I_{s,\ell},I_{e,\ell},I_{m,\ell}\right\}$, and the entries are sampled from the corresponding set of distributions $\mathbb{p}_{\mathcal{R}}=\{\mathbb{p}_{\epsilon_{\mu}},\mathbb{p}_{\epsilon_{\sigma}},\mathbb{p}_{\lambda},\mathfrak{S}_{\mathcal{I}_{s}},\mathfrak{S}_{\mathcal{I}_{e}},\mathfrak{S}_{\mathcal{B}}\}$. $\mathbb{p}_{D_{i}}$ denotes the probability distribution of source data in device $i\in\mathcal{M}$, and $\mathfrak{S}_{\mathcal{I}_{s}}$ and $\mathfrak{S}_{\mathcal{I}_{e}}$ are supersets of all unique permutations of the sets $\mathcal{I}_{s}$ and $\mathcal{I}_{e}$, respectively. In turn, $\mathcal{I}_{s}$ and $\mathcal{I}_{e}$ are binary vectors of size $B$, with the number of $1$s being equal to $B_{s}$ and $B_{e}$, respectively. For a given layer $\ell$, each device $i$ utilizes the received styles of a neighboring device $j$ in the training process, i.e., $\psi_{j_{\ell}}=\begin{bmatrix}\bar{\mu}_{j_{\ell}}&\bar{\sigma}_{j_{\ell}}&\Sigma_{\mu,j_{\ell}}^{2}&\Sigma_{\sigma,j_{\ell}}^{2}\end{bmatrix}^{T}$ as the $j$th element of $\Psi$ in Eq. (20). In essence, StyleDDG is applying a style exploration strategy to the outputs of the model’s layer $\ell$ with probability $p_{\ell}$, so that the overall model is modified as
@@
in which $\tilde{h}_{\theta_{\ell},\psi_{j_{\ell}},\mathcal{R}_{\ell}}=h_{\theta_{\ell}}$ if $\ell\notin\pi$. If $\ell\in\pi$, we define $x_{s}=[x_{k}]_{k\in\{1,...,B\};(I_{s,\ell})_{k}=1}^{T}$ as a subset of the mini-batch $x$ based on the indices $I_{s,\ell}$ that we will shift to the styles of a neighboring device, and $x_{s}^{c}=[x_{k}]_{k\in\{1,...,B\};(I_{s,\ell})_{k}=0}^{T}$ as the ones which will not be shifted. Then the $\operatorname{StyleExplore}$ update that occurs at that layer can be expressed as
@@
where we first apply the $\operatorname{StyleShift}$ function on a given mini-batch $x$ to randomly shift $B_{s}$ of the points to new styles received from a neighbor, and then the $\operatorname{StyleExplore}$ layer takes the outputs of the $\operatorname{StyleShift}$ function to randomly extrapolate $B_{e}$ of those styles outwards for style exploration.
@@ V Convergence Analysis
In this section, we characterize the convergence behavior of StyleDDG. Papers such as [43, 41, 44] prove that in decentralized optimization of general non-convex models, convergence to stationary points can be guaranteed if the local loss functions are smooth, i.e., their gradients are Lipschitz continuous. To this end, we will discuss the conditions under which the smoothness of local loss functions in StyleDDG given in Eq. (20) can be guaranteed as well. In other words, our goal is to prove that the upper bound $\|\nabla{F}_{i}(\theta,\Psi)-\nabla{F}_{i}(\theta^{\prime},\Psi^{\prime})\|\leq\beta_{i}\|\theta-\theta^{\prime}\|$ can be established for each device in StyleDDG, where $\beta_{i}\in\mathbb{R}^{+}$ is the smoothness coefficient.
@@ V-A Assumptions
Assumption V.1 is necessary to ensure the convergence of gradient-based methods, and it indicates that the local gradients do not change arbitrarily quickly [45]. The conventional smoothness assumption made in [45, 43] is of the form $\|\nabla{F}_{i}(\theta)-\nabla{F}_{i}(\theta^{\prime})\|\leq\beta_{\theta,i}\|\theta-\theta^{\prime}\|$. These works do not modify their CNN models with any style-based statistics, and the local loss functions are parameterized by $\theta$ alone. Assumption V.1 extends this to incorporate smoothness in terms of style vectors as well, i.e., the local loss functions do not change arbitrarily quickly with the style statistics.
@@
which is the standard smoothness assumption for a univariate function with $\beta_{\phi,i}\in\mathbb{R}^{+}$. Note that Eq. (24) is stricter than the multivariate smoothness assumption we use in Eq. (23) since we can obtain the former from the latter using triangle inequality.
@@
    (Liptschitz continuity of layer outputs) We have $\|h_{\theta_{\ell}}(x_{i})_{b,c,h,w}-h_{\theta_{\ell}^{\prime}}(x_{i})_{b,c,h,w}\|\leq D_{i,\ell}\|\theta-\theta^{\prime}\|,$ where $\theta,\theta^{\prime}\in\mathbb{R}^{n}$ are two arbitrary model parameter vectors and $D_{i,\ell}\in\mathbb{R}^{+}$ is a positive scalar. Also, we define $D_{i}=\max_{\ell}\{D_{i,\ell}\}$ and $D=\max_{i}\{D_{i}\}$.
@@
    (Bounded layer outputs) For all $\theta\in\mathbb{R}^{n}$, there exists $U_{i,\ell}\in\mathbb{R}^{+}$ such that $\|h_{\theta_{\ell}}(x_{i})_{b,c,h,w}\|\leq U_{i,\ell}<\infty.$ Also, we define $U_{i}=\max_{\ell}\{U_{i,\ell}\}$ and $U=\max_{i}\{U_{i}\}$.
@@ V-B Main Results
(Lipschitz continuity of style statistics) Let Assumption V.3 hold. Then the style statistics from Eq. (15) are Lipschitz continuous at each layer/block $\ell=\{1,...,L\}$ of the CNN $\theta$ for all devices $i\in\mathcal{M}$, and for all channels $c\in\{1,...,C\}$. Specifically:
@@
    $\|\bar{\mu}_{\theta,\ell}(x_{i})_{c}-\bar{\mu}_{\theta^{\prime},\ell}(x_{i})_{c}\|\leq D_{i,\ell}\|\theta-\theta^{\prime}\|$
@@
    $\|\bar{\sigma}_{\theta,\ell}(x_{i})_{c}-\bar{\sigma}_{\theta^{\prime},\ell}(x_{i})_{c}\|\leq\frac{4U_{i}D_{i,\ell}}{\sqrt{\eta}}\|\theta-\theta^{\prime}\|$
@@
    $\|\Sigma_{\mu,\theta,\ell}^{2}(x_{i})_{c}-\Sigma_{\mu,\theta^{\prime},\ell}^{2}(x_{i})_{c}\|\leq 4U_{i}D_{i,\ell}\|\theta-\theta^{\prime}\|$
@@
    $\|\Sigma_{\sigma,\theta,\ell}^{2}(x_{i})_{c}\!-\!\Sigma_{\sigma,\theta^{\prime},\ell}^{2}(x_{i})_{c}\|\!\leq\!4U_{i}D_{i,\ell}\big{(}1\!+\!\frac{2\sqrt{2}U_{i}}{\sqrt{\eta}}\big{)}\|\theta\!-\!\theta^{\prime}\|$
Sketch of Proof. [^1] Due to space limitations, we are only able to provide a sketch of each proof. In the camera-ready, we will provide a technical report with full proofs. For each of the style statistics, we start the derivation of the upper bound using their corresponding definitions in Eq. (15). Part (a) requires only the Lipschitzness of CNN outputs $h_{\theta}(x)$ (Assumption V.3-(a)). Part (b) requires us to establish the Lipschitzness of $h_{\theta}^{2}(x)$ (where the boundedness of CNN outputs defined in Assumption V.3-(b) is used), and then use the lower bound for $\sigma_{\theta}(x)$ enforced by the numerical stability value $\eta$ to conclude the proof. A crucial piece of part (b)’s proof is to derive the bounds in such a way that the stability parameter $\eta$ can be used to obtain the bounds, which has not been investigated in earlier works [28]. Finally, parts (c) and (d) use parts (a) and (b) to obtain upper bounds for their variances, respectively.
@@
where $\beta_{i}=\beta_{\theta,i}+mL(1+4U(\frac{1}{\sqrt{\eta}}+2+\frac{2\sqrt{2}U}{\sqrt{\eta}}))D\beta_{\Psi,i}$, $\beta_{\theta,i}$ and $\beta_{\Psi,i}$ are provided in Assumption V.1, $m$ is the total number of devices, $L$ is the number of layers which StyleDDG is applied to, $U$ and $D$ are the bounds from Assumption V.3, and $\eta>0$ is the small numerical stability constant.
Sketch of Proof. With the Lipschitzness of all four individual style statistics of each device and layer/block established in Proposition V.4, we first establish the Lipschitz continuity of their concatenation $\psi_{\theta,\ell}(x_{i})=\begin{bmatrix}\bar{\mu}_{\theta,\ell}(x_{i})&\bar{\sigma}_{\theta,\ell}(x_{i})&\Sigma_{\mu,\theta,\ell}^{2}(x_{i})&\Sigma_{\sigma,\theta,\ell}^{2}(x_{i})\end{bmatrix}^{T}$. Next, we focus on the Lipschitzness of style statistics across all the layers of the CNN for a given client $i\in\mathcal{M}$, i.e., $\psi_{\theta}(x_{i})=\begin{bmatrix}\psi_{\theta,1}(x_{i})&\cdots&\psi_{\theta,L}(x_{i})\end{bmatrix}^{T}$. This helps us to ultimately prove the Lipschitzness of full style statistics matrix $\Psi_{\theta}(X)=\begin{bmatrix}\psi_{\theta}(x_{1})&\cdots&\psi_{\theta}(x_{m})\end{bmatrix}^{T}$.
@@
Let Assumptions V.1 and V.3 hold, and a constant learning rate $\alpha^{(k)}=\alpha^{(0)}/{\sqrt{K+1}}$ be employed, where $K$ is the maximum number of iterations that StyleDDG will run. Then, for general non-convex loss model, we have
@@
where $\beta=(1/m)\sum_{i=1}^{m}{\beta_{i}}$, $F^{\star}=\min_{\theta\in\mathbb{R}^{n}}\{F(\theta)\}$, $m$ denotes the number of devices, $\bar{\theta}^{(k)}=(1/m)\sum_{i=1}^{m}{\theta_{i}^{(k)}}$ is the average model among all devices at iteration $k$, $\rho$ is the spectral radius of the connected network graph $\mathcal{G}$, and $\beta$ is the smoothness parameter given by Proposition V.5.
Sketch of Proof. With the univariate smoothness of local loss functions established in Proposition V.5, we can build upon established theoretical results in decentralized optimization to provide convergence guarantees. The proof can follow steps similar to e.g., Theorem 4.12 in [41] or Lemma 16 in [43].
Theorem V.6 shows that StyleDDG reaches a stationary point for general non-convex models at a rate of $\mathcal{O}(1/\sqrt{K})$. To be specific, after running StyleDDG for $K$ iterations with a learning rate proportional to $1/\sqrt{K+1}$, the running average ($\frac{1}{K+1}\sum_{k=0}^{K}{(\cdot)}$) of global gradient norm expectations ($\mathbb{E}{[\|\nabla{F}(\cdot)\|]}$), evaluated at the average model $\bar{\theta}^{(k)}$, is proportional to $1/\sqrt{K+1}$. This bound becomes arbitrarily small as the training progresses, i.e., by increasing $K$. Further, in Theorem V.6, learning model-related parameters of Assumptions V.1 and V.3, i.e., multivariate smoothness parameters $\beta_{\theta,i}$ and $\beta_{\Psi,i}$, Lipschitz constant $D_{i,\ell}$ and upper bound $U_{i,\ell}$, are all captured in the univariate smoothness parameter $\beta$ derived in Proposition V.5. We see in Eq. (25) that a lower smoothness value $\beta$, as well as a better-connected D2D graph (i.e., smaller spectral radius $\rho$), results in a smaller upper bound in Eq. (25), accelerating the convergence speed accordingly.

### v4

@@ IV-B1 Global Loss
where $F_{i}(\theta,\Psi_{\theta,i}(X))$ is the local loss function of each device $i\in\mathcal{M}$ evaluated with the model $\theta$ and the combined batch $X$ over devices, in which $\Psi_{\theta,i}(X)\subset\Psi_{\theta}(X)$ is the set of styles belonging to the neighbors of device $i$. Note that in Eq. (19), if instead of $\Psi_{\theta,i}(X)$ we provide the full style information of $\Psi_{\theta}(X)$ to each individual device, we recover the loss function for DG in star-topology FL, as considered in StableFDG [park2024stablefdg].
@@ IV-B2 Local Loss
$F_{i}(\theta,\Psi)\hskip-2.84526pt=\hskip-11.38109pt\sum_{\pi\in\mathcal{P}(L)}{\hskip-5.69054pt\left(\prod_{\ell\in\pi}{p_{\ell}}\hskip-2.84526pt\right)\hskip-2.84526pt\mathbb{E}_{\hskip-8.53581pt\begin{subarray}{c}(x,y)\sim\mathbb{p}_{D_{i}}\\ \{j_{\ell}\in\mathcal{M}\setminus\{i\}\}_{\ell\in\pi}\\ \{\mathcal{R}_{\ell}\sim\mathbb{p}_{\mathcal{R}}\}_{\ell\in\pi}\end{subarray}\hskip-11.38109pt}{\left[\mathscr{L}{(\tilde{h}_{\theta,\{\psi_{j_{\ell}},\mathcal{R}_{\ell}\}_{l\in\pi}}(x),y)}\right]}}\hskip-1.42262pt,\vskip-5.69054pt$ (20)
in which the expectation is taken over the set of random variables $\mathcal{R}_{\ell}=\left\{\epsilon_{\mu,\ell},\epsilon_{\sigma,\ell},\lambda_{\ell},I_{s,\ell},I_{e,\ell},I_{m,\ell}\right\}$, and the entries are sampled from the corresponding set of distributions $\mathbb{p}_{\mathcal{R}}=\{\mathbb{p}_{\epsilon_{\mu}},\mathbb{p}_{\epsilon_{\sigma}},\mathbb{p}_{\lambda},\mathfrak{S}_{\mathcal{I}_{s}},\mathfrak{S}_{\mathcal{I}_{e}},\mathfrak{S}_{\mathcal{B}}\}$. $\mathbb{p}_{D_{i}}$ denotes the probability distribution of source data in device $i\in\mathcal{M}$, and $\mathfrak{S}_{\mathcal{I}_{s}}$ and $\mathfrak{S}_{\mathcal{I}_{e}}$ are supersets of all unique permutations of the sets $\mathcal{I}_{s}$ and $\mathcal{I}_{e}$, respectively. In turn, $\mathcal{I}_{s}$ and $\mathcal{I}_{e}$ are binary vectors of size $B$, with the number of $1$’s equal to $B_{s}$ and $B_{e}$, respectively. For a given layer $\ell$, each device $i$ uses the received styles of a neighboring device $j$ in the training process, i.e., $\psi_{j_{\ell}}=\begin{bmatrix}\bar{\mu}_{j_{\ell}}&\bar{\sigma}_{j_{\ell}}&\Sigma_{\mu,j_{\ell}}^{2}&\Sigma_{\sigma,j_{\ell}}^{2}\end{bmatrix}^{T}$ as the $j$th element of $\Psi$ in Eq. (20). In essence, StyleDDG applies a style exploration strategy to the output of the model’s layer $\ell$ with probability $p_{\ell}$, so that the overall model is modified as
@@
in which $\tilde{h}_{\theta_{\ell},\psi_{j_{\ell}},\mathcal{R}_{\ell}}=h_{\theta_{\ell}}$ if $\ell\notin\pi$. If $\ell\in\pi$, we define $x_{s}=[x_{k}]_{k\in\{1,...,B\};(I_{s,\ell})_{k}=1}^{T}$ as a subset of the mini-batch $x$ based on the indices $I_{s,\ell}$ that we will shift to the styles of a neighboring device and $x_{s}^{c}=[x_{k}]_{k\in\{1,...,B\};(I_{s,\ell})_{k}=0}^{T}$ as the ones which will not be shifted. Then the $\operatorname{StyleExplore}$ update that occurs in that layer can be expressed as
@@
where we first apply the $\operatorname{StyleShift}$ function on a given mini-batch $x$ to randomly shift $B_{s}$ of the points to new styles received from a neighbor, and then the $\operatorname{StyleExplore}$ layer takes the output of the $\operatorname{StyleShift}$ function to randomly extrapolate $B_{e}$ of those styles outwards for style exploration.
@@ V Convergence Analysis
In this section, we characterize the convergence behavior of StyleDDG. Papers such as [koloskova2020unified, zehtabi2024decentralized, xin2021improved] prove that in decentralized optimization of general non-convex models, convergence to stationary points can be guaranteed if the local loss functions are smooth, i.e., their gradients are Lipschitz continuous. To this end, we will discuss the conditions under which the smoothness of local loss functions in StyleDDG given in Eq. (20) can be guaranteed as well. In other words, our goal is to prove that the upper bound $\|\nabla{F}_{i}(\theta,\Psi)-\nabla{F}_{i}(\theta^{\prime},\Psi^{\prime})\|\leq\beta_{i}\|\theta-\theta^{\prime}\|$ can be established for each device in StyleDDG, where $\beta_{i}\in\mathbb{R}^{+}$ is the smoothness coefficient.
@@ V-A Assumptions
The Assumption V.1 is necessary to ensure the convergence of gradient-based methods, and it ensures that the local gradients do not change arbitrarily fast [bottou2018optimization]. The conventional smoothness assumption made in [bottou2018optimization, koloskova2020unified] is of the form $\|\nabla{F}_{i}(\theta)-\nabla{F}_{i}(\theta^{\prime})\|\leq\beta_{\theta,i}\|\theta-\theta^{\prime}\|$. These works do not modify their CNN models with any style-based statistics, and the local loss functions are parameterized by $\theta$ alone. Assumption V.1 extends this to incorporate smoothness in terms of style vectors as well, i.e., the local loss functions do not change arbitrarily quickly with the style statistics.
@@
which is the standard smoothness assumption for a univariate function with $\beta_{\phi,i}\in\mathbb{R}^{+}$. Note that Eq. (24) is stricter than the multivariate smoothness assumption in Eq. (23) since we can obtain the former from the latter using the triangle inequality.
@@
    (Liptschitz continuity of layer outputs) We have $\|h_{\theta_{\ell}}(x_{i})_{b,c,h,w}-h_{\theta_{\ell}^{\prime}}(x_{i})_{b,c,h,w}\|\leq D_{i,\ell}\|\theta-\theta^{\prime}\|,$ where $\theta,\theta^{\prime}\in\mathbb{R}^{n}$ are any two model parameter vectors and $D_{i,\ell}\in\mathbb{R}^{+}$ is a positive scalar. We also define $D_{i}=\max_{\ell}\{D_{i,\ell}\}$ and $D=\max_{i}\{D_{i}\}$.
@@
    (Bounded layer output) For all $\theta\in\mathbb{R}^{n}$, there exists $U_{i,\ell}\in\mathbb{R}^{+}$ such that $\|h_{\theta_{\ell}}(x_{i})_{b,c,h,w}\|\leq U_{i,\ell}<\infty$. We also define $U_{i}=\max_{\ell}\{U_{i,\ell}\}$ and $U=\max_{i}\{U_{i}\}$.
@@ V-B Main Results
(Lipschitz continuity of style statistics) Let Assumption V.3 hold. Then the style statistics from Eq. (15) are Lipschitz continuous at each layer/block $\ell=\{1,...,L\}$ of CNN $\theta$ for all devices $i\in\mathcal{M}$, and for all channels $c\in\{1,...,C\}$. Specifically:
@@
    $\|\bar{\mu}_{\theta,\ell}(x_{i})_{c}-\bar{\mu}_{\theta^{\prime},\ell}(x_{i})_{c}\|\leq D_{i,\ell}\|\theta-\theta^{\prime}\|$.
@@
    $\|\bar{\sigma}_{\theta,\ell}(x_{i})_{c}-\bar{\sigma}_{\theta^{\prime},\ell}(x_{i})_{c}\|\leq\frac{4U_{i}D_{i,\ell}}{\sqrt{\eta}}\|\theta-\theta^{\prime}\|$.
@@
    $\|\Sigma_{\mu,\theta,\ell}^{2}(x_{i})_{c}-\Sigma_{\mu,\theta^{\prime},\ell}^{2}(x_{i})_{c}\|\leq 4U_{i}D_{i,\ell}\|\theta-\theta^{\prime}\|$.
@@
    $\|\Sigma_{\sigma,\theta,\ell}^{2}(x_{i})_{c}\!-\!\Sigma_{\sigma,\theta^{\prime},\ell}^{2}(x_{i})_{c}\|\!\leq\!4U_{i}D_{i,\ell}\big(1\!+\!\frac{2\sqrt{2}U_{i}}{\sqrt{\eta}}\big)\|\theta\!-\!\theta^{\prime}\|$.
Proof. See our technical report [zehtabi2025decentralized].
@@
where $\beta_{i}=\beta_{\theta,i}+mL(1+4U(\frac{1}{\sqrt{\eta}}+2+\frac{2\sqrt{2}U}{\sqrt{\eta}}))D\beta_{\Psi,i}$, $\beta_{\theta,i}$ and $\beta_{\Psi,i}$ are provided in Assumption V.1, $m$ is the total number of devices, $L$ is the number of layers that StyleDDG is applied to, $U$ and $D$ are the bounds of Assumption V.3, and $\eta>0$ is the small numerical stability constant.
Proof. See our technical report [zehtabi2025decentralized].
@@
Let Assumptions V.1 and V.3 hold, and a constant learning rate $\alpha^{(k)}=\alpha^{(0)}/{\sqrt{K+1}}$ be employed where $K$ is the maximum number of iterations that StyleDDG will run for. Then, for a general non-convex loss function, we have
@@
where $\beta=(1/m)\sum_{i=1}^{m}{\beta_{i}}$, $F^{\star}=\min_{\theta\in\mathbb{R}^{n}}\{F(\theta)\}$, $m$ denotes the number of devices, $\bar{\theta}^{(k)}=(1/m)\sum_{i=1}^{m}{\theta_{i}^{(k)}}$ is the average model among all devices in iteration $k$, $\rho$ is the spectral radius of the connected network graph $\mathcal{G}$, and $\beta$ is the smoothness parameter given by Proposition V.5.
Proof. See our technical report [zehtabi2025decentralized].
Theorem V.6 shows that StyleDDG reaches a stationary point for general non-convex models at a rate of $\mathcal{O}(1/\sqrt{K})$. To be specific, after running StyleDDG for $K$ iterations with a learning rate proportional to $1/\sqrt{K+1}$, the running average ($\frac{1}{K+1}\sum_{k=0}^{K}{(\cdot)}$) of global gradient norm expectations ($\mathbb{E}{[\|\nabla{F}(\cdot)\|]}$), evaluated at the average model $\bar{\theta}^{(k)}$, is proportional to $1/\sqrt{K+1}$. This bound becomes arbitrarily small as training progresses, i.e., as $K$ increases. Additionally, in Theorem V.6, ML model-related parameters of Assumptions V.1 and V.3, i.e., multivariate smoothness parameters $\beta_{\theta,i}$ and $\beta_{\Psi,i}$, Lipschitz constant $D_{i,\ell}$ and the upper bound $U_{i,\ell}$, are all captured in the univariate smoothness parameter $\beta$ derived in Proposition V.5. We see in Eq. (25) that a lower smoothness value $\beta$, as well as a better-connected D2D graph (i.e., smaller spectral radius $\rho$), results in a smaller upper bound in Eq. (25), accelerating the convergence speed accordingly.